# Almost all 9-regular graphs have a modulo-5 orientation

### v4

@@ 1. Introduction
To orient a graph $G=(V,E)$, convert each edge $e=\{x,y\}$ into either the ordered pair $e=(x,y)$ or $e=(y,x)$, to represent the edge being oriented from $x$ to $y$ or from $y$ to $x$, respectively. We denote the set of all edges oriented towards $v$ as $D^{+}(v)$ and the set of all edges oriented away from $v$ as $D^{-}(v)$; a $k$-flow is said to be a nowhere-zero $k$-flow if $f(e)\neq 0$ for all $e\in E$. We define the in-degree of $v$ as $d^{+}(v):=|D^{+}(v)|$ and the out-degree of $v$ as $d^{-}(v):=|D^{-}(v)|$. In 1972 Tutte famously conjectured the following (for instance, see Bondy [5, Open Problem 48] or Jensen [12, Section 13.3]):
@@
Though Tutte’s nowhere-zero 3-flow conjecture has attracted considerable attention, it has yet to be proven. For a significant period of time it was not known whether or not there exists a fixed constant $k$ such that every $k$-edge connected graph has a nowhere-zero 3-flow (this is known as the weak 3-flow conjecture of Jaeger). It was proved for $k\geq c\log_{2}n$ for $n$-vertex graphs by Alon, Linial, and Meshulam in 1991 [1], as well as by Lai and Zhang in 1992 [13]. The weak $3$-flow conjecture was first proved in 2012 by Thomassen [18], who showed that every $8$-edge-connected graph admits a nowhere-zero 3-flow. This was improved to $k=6$ in 2011 by Lovász, Thomassen, Wu, and Zhang [14]. It is a well-known fact that a graph admits a nowhere-zero 3-flow if and only if it has a nowhere-zero 3-flow over $\mathbb{Z}_{3}$, for instance see Seymour [17] for more discussion. The following is a reformulation of Tutte’s nowhere-zero 3-flow conjecture ([12, Section 13.3], [5, Open Problem 48]):
@@ 2. Overview of Our Proof of the Main Result
$\frac{\mathbb{E}(Y[X_{1}]_{\ell_{1}}\cdots[X_{j}]_{\ell_{j}})}{\mathbb{E}Y}\sim\prod\limits_{i=1}^{j}\mu_{i}^{j_{i}}=\prod\limits_{i=1}^{j}\left(\lambda_{i}(1+\delta_{i})\right)^{j_{i}}=\prod\limits_{i=1}^{j}\left(\frac{8^{i}}{2i}\left(1-\left(\frac{2}{9}\right)^{i}\right)\right)^{j_{i}}$
@@
Recall that $\lambda_{k}\frac{8^{k}}{2k}$ and $\delta_{k}=-\left(\frac{2}{9}\right)^{k}$. Using the fact that $-\log(1-x)=\sum_{k\geq 1}x_{k}/k!$,
$\exp\left(\sum\limits_{k\geq 1}\lambda_{k}\delta_{k}^{2}\right)=\sum\limits_{k\geq 1}\frac{8^{k}}{2k}\left(\frac{4}{81}\right)^{k}=\exp\left(\frac{1}{2}-\log\left(1-\frac{2^{5}}{3^{4}}\right)\right)=\frac{9}{7}.$
@@ 4.1. Multivariate Calculus
where $\textbf{z}=\textbf{z}(\textbf{k})=\textbf{k}/n$. The product within the sum is merely the result of applying Stirling’s approximation $(s!\sim\sqrt{2\pi n}(s/e)^{s})$. The error term, the polynomial component, and the exponential component are represented by $r(\textbf{z})$, $g(\textbf{z})$, and $\exp(nf(\textbf{z}))$ respectively and defined below. Note that since $\sqrt{2\pi n}(\frac{n}{e})^{n}e^{\left(\frac{1}{12n}-\frac{1}{360n^{3}}\right)}<n<\sqrt{2\pi n}(\frac{n}{e})^{n}e^{\left(\frac{1}{12n}\right)}$, $r(\textbf{z})=O(1)$ for all z and $r(\textbf{z})\sim 1$ if z is bounded away from the boundary of
@@
We can use the fact that $z_{211}z_{100}=z_{200}z_{111}$ and $\mathcal{R}(P_{211},P_{200},z_{101})$ to demonstrate that $z_{211}=z_{200}$, which then implies that $z_{111}=z_{100}$.
@@
We can use the fact that $z_{210}z_{101}=z_{201}z_{110}$ and $\mathcal{R}(P_{210},P_{201},z_{100})$ to demonstrate that $z_{210}=z_{201}$, which then implies that $z_{110}=z_{101}$.
@@
$z=\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}.$
@@
$z=\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2},$
@@
$c_{4}=\frac{1/2-\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}}{21\left(\frac{b}{9/2-b}\right)^{2}+14\left(\frac{b}{9/2-b}\right)+1}.$
@@
$c_{3}=14\left(\frac{b}{9/2-b}\right)\frac{1/2-\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}}{21\left(\frac{b}{9/2-b}\right)^{2}+14\left(\frac{b}{9/2-b}\right)+1}.$
@@
$0<c_{2}=c_{2}^{*}(b)<\frac{1/2}{\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)}.$
@@
$c<\frac{1/2}{\left(\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)}$
@@ 4.2. Analyzing the Boundary Points
Setting $\frac{\delta f}{\delta z_{110}}=\frac{\delta f}{\delta z_{101}}$ implies
@@
Setting $\frac{\delta f}{\delta z_{210}}=\frac{\delta f}{\delta z_{201}}$ implies
@@
Together these imply that $z_{211}z_{100}=z_{200}z_{111}$. Setting $\frac{\delta f}{\delta z_{111}}=\frac{\delta f}{\delta z_{100}}$ implies that
@@
which in turn implies that $z_{111}=z_{100}$. Setting $\frac{\delta f}{\delta z_{211}}=\frac{\delta f}{\delta z_{200}}$ implies that
@@
Since $r$ and $g$ are polynomially bounded it follows that the terms in the summation (6) for which $\textbf{k}/n\in J/J_{0}$ are bounded by $(81/8)^{n}e^{-\Omega(n^{1/5})}$. Since there are polynomially many such terms their sum is $(81/8)^{n}e^{-\Omega(n/5)}$.
@@ 5. Joint Factorial Moments
There are $\frac{(9\cdot 8)^{k}}{2k}\cdot\frac{n!}{(n-k)!}$ different ways to choose vertices that make up the cycle. There are $i$ vertices with in-degree 2, $b$ of which are in-vertices. Since the sum of in-degrees in the cycle must equal the number of out-degrees in the cycle, there are $i$ vertices in the cycle with out-degree 2, $c$ of which are out-vertices. This implies that $k-2i$ vertices in the cycle have an in-degree and out-degree of 1 in the context of the cycle. We let $a_{i}=2{k\choose 2i}$ refer to the number of orientations. The number of choices of in-vertices among $i$-vertices with in-degree 2 is (${i\choose b}$. Likewise, there are (${i\choose c}$ possible choices of out-vertices among $i$ vertices with out-degree 2. Turning our attention to the vertices outside of the cycle, we see that there are (${n-2i\choose n/2-i-b+c}$ ways to pair the in- and out-vertices that are not in $C$.
We now consider the distribution of special points between our orientations. There are (${7\choose 2}^{i-b}$ ways to select special points in vertices that have an in-degree of 2 in the cycle and are out-vertices in the graph. Likewise, there are (${7\choose 2}^{i-c}$ ways to select special points in vertices that have an out-degree of 2 in the cycle and which are in-vertices in the graph. The special points for vertices with in-degree 2 in $C$ and which are in-vertices and for vertices with out-degree 2 in $C$ and which are out-vertices do not need to be identified. There are $7^{k-2i}$ ways to select special points among vertices in $C$ which have an in-degree of 1 and an out-degree of 1. There are (${9\choose 2}^{n-k}$ ways to select special points for vertices outside of $C$. Given that $9n/2-k$ must be in-vertices, there are $(9n/2-k)!$ different ways to pair in-vertices with out-vertices. We are now ready to verify Lemma 2.5.
@@
We calculate $\frac{\mathbb{E}[X_{k}Y]}{\mathbb{E}[Y]}$ as follows:
@@
Since $n\to\infty$ and $k$ is finite Sterling’s formula implies that $\frac{(n-2i)!}{n!}\sim\left(\frac{n}{e}\right)^{-2i}$ and
@@
Returning to our calculation of $\frac{\mathbb{E}[X_{k}Y]}{\mathbb{E}Y}$ gives us
@@
Since Corollary 2.19 from [4] stipulates that the cycles are asymptotically independent, we are justified in further concluding that, for every finite sequence $j_{1},j_{2},\ldots,j_{k}$ of non-negative integers, the following holds:
@@
where $\delta_{i}=\frac{\mu_{i}}{\lambda_{i}}-1$ to satisfy condition (c). For the $k$ that we concerned ourselves with above $\delta_{k}=-\left(\frac{2}{9}\right)^{k}$.

### v5

@@ 1. Introduction
To orient a graph $G=(V,E)$, convert each edge $e=\{x,y\}$ into either the ordered pair $e=(x,y)$ or $e=(y,x)$, to represent the edge being oriented from $x$ to $y$ or from $y$ to $x$, respectively. We denote the set of all edges oriented towards $v$ as $D^{+}(v)$ and the set of all edges oriented away from $v$ as $D^{-}(v)$; a $k$-flow is said to be a nowhere-zero $k$-flow if $f(e)\neq 0$ for all $e\in E$. We define the in-degree of $v$ as $d^{+}(v):=|D^{+}(v)|$ and the out-degree of $v$ as $d^{-}(v):=|D^{-}(v)|$. In 1972 Tutte famously conjectured the following (for instance, see Bondy [5, Open Problem 48] or Jensen [13, Section 13.3]):
@@
Though Tutte’s nowhere-zero 3-flow conjecture has attracted considerable attention, it has yet to be proven. For a significant period of time it was not known whether or not there exists a fixed constant $k$ such that every $k$-edge connected graph has a nowhere-zero 3-flow (this is known as the weak 3-flow conjecture of Jaeger). It was proved for $k\geq c\log_{2}n$ for $n$-vertex graphs by Alon, Linial, and Meshulam in 1991 [1], as well as by Lai and Zhang in 1992 [14]. The weak $3$-flow conjecture was first proved in 2012 by Thomassen [19], who showed that every $8$-edge-connected graph admits a nowhere-zero 3-flow. This was improved to $k=6$ in 2011 by Lovász, Thomassen, Wu, and Zhang [15]. It is a well-known fact that a graph admits a nowhere-zero 3-flow if and only if it has a nowhere-zero 3-flow over $\mathbb{Z}_{3}$, for instance see Seymour [18] for more discussion. The following is a reformulation of Tutte’s nowhere-zero 3-flow conjecture ([13, Section 13.3], [5, Open Problem 48]):
@@ 2. Overview of Our Proof of the Main Result
$\frac{\mathbb{E}(Y[X_{1}]_{\ell_{1}}\cdots[X_{j}]_{\ell_{j}})}{\mathbb{E}Y}\sim\prod\limits_{i=1}^{j}\mu_{i}^{\ell_{i}}=\prod\limits_{i=1}^{j}\left(\lambda_{i}(1+\delta_{i})\right)^{\ell_{i}}=\prod\limits_{i=1}^{j}\left(\frac{8^{i}}{2i}\left(1-\left(\frac{2}{9}\right)^{i}\right)\right)^{\ell_{i}}$
@@
Recall that $\lambda_{k}=\frac{8^{k}}{2k}$ and $\delta_{k}=-\left(\frac{2}{9}\right)^{k}$. Using the fact that $-\log(1-x)=\sum_{k\geq 1}x^{k}/k$,
$\exp\left(\sum\limits_{k\geq 1}\lambda_{k}\delta_{k}^{2}\right)=\exp\left(\sum\limits_{k\geq 1}\frac{8^{k}}{2k}\left(\frac{4}{81}\right)^{k}\right)=\exp\left(\frac{1}{2}-\log\left(1-\frac{2^{5}}{3^{4}}\right)\right)=\frac{9}{7}.$
@@ 4.1. Multivariate Calculus
where $\textbf{z}=\textbf{z}(\textbf{k})=\textbf{k}/n$. The product within the sum is merely the result of applying Stirling’s approximation $(s!\sim\sqrt{2\pi n}(s/e)^{s})$. The error term, the polynomial component, and the exponential component are represented by $r(\textbf{z})$, $g(\textbf{z})$, and $\exp(nf(\textbf{z}))$ respectively and defined below. Note that since $\sqrt{2\pi n}(\frac{n}{e})^{n}e^{\left(\frac{1}{12n}-\frac{1}{360n^{3}}\right)}<n!<\sqrt{2\pi n}(\frac{n}{e})^{n}e^{\left(\frac{1}{12n}\right)}$, $r(\textbf{z})=O(1)$ for all z and $r(\textbf{z})\sim 1$ if z is bounded away from the boundary of
@@
We can use the fact that $z_{211}z_{100}=z_{200}z_{111}$ and $\mathcal{R}(P_{211},P_{200},z_{101})=0$ to demonstrate that $z_{211}=z_{200}$, which then implies that $z_{111}=z_{100}$.
@@
We can use the fact that $z_{210}z_{101}=z_{201}z_{110}$ and $\mathcal{R}(P_{210},P_{201},z_{100})=0$ to demonstrate that $z_{210}=z_{201}$, which then implies that $z_{110}=z_{101}$.
@@
$z=\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}.$
@@
$z=\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2},$
@@
$c_{4}=\frac{1/2-\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}}{21\left(\frac{b}{9/2-b}\right)^{2}+14\left(\frac{b}{9/2-b}\right)+1}.$
@@
$c_{3}=14\left(\frac{b}{9/2-b}\right)\frac{1/2-\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)c_{2}}{21\left(\frac{b}{9/2-b}\right)^{2}+14\left(\frac{b}{9/2-b}\right)+1}.$
@@
$0<c_{2}=c_{2}^{*}(b)<\frac{1/2}{\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)}.$
@@
$c<\frac{1/2}{\left(21\left(\frac{9/2-b}{b}\right)^{2}+14\left(\frac{9/2-b}{b}\right)+1\right)}$
@@ 4.2. Analyzing the Boundary Points
Setting $\frac{\delta\bar{f}}{\delta z_{110}}=\frac{\delta\bar{f}}{\delta z_{101}}$ implies
@@
Setting $\frac{\delta\bar{f}}{\delta z_{210}}=\frac{\delta\bar{f}}{\delta z_{201}}$ implies
@@
Together these imply that $z_{211}z_{100}=z_{200}z_{111}$. Setting $\frac{\delta\bar{f}}{\delta z_{111}}=\frac{\delta\bar{f}}{\delta z_{100}}$ implies that
@@
which in turn implies that $z_{111}=z_{100}$. Setting $\frac{\delta\bar{f}}{\delta z_{211}}=\frac{\delta\bar{f}}{\delta z_{200}}$ implies that
@@
Since $r$ and $g$ are polynomially bounded it follows that the terms in the summation (6) for which $\textbf{k}/n\in J/J_{0}$ are bounded by $(81/8)^{n}e^{-\Omega(n^{1/5})}$. Since there are polynomially many such terms their sum is $(81/8)^{n}e^{-\Omega(n^{1/5})}$.
@@ 5. Joint Factorial Moments
There are $\frac{(9\cdot 8)^{k}}{2k}\cdot\frac{n!}{(n-k)!}$ different ways to choose vertices that make up the cycle. There are $i$ vertices with in-degree 2, $b$ of which are in-vertices. Since the sum of in-degrees in the cycle must equal the number of out-degrees in the cycle, there are $i$ vertices in the cycle with out-degree 2, $c$ of which are out-vertices. This implies that $k-2i$ vertices in the cycle have an in-degree and out-degree of 1 in the context of the cycle. We let $a_{i}=2{k\choose 2i}$ refer to the number of orientations. The number of choices of in-vertices among $i$-vertices with in-degree 2 is (${i\choose b}$. Likewise, there are (${i\choose c}$ possible choices of out-vertices among $i$ vertices with out-degree 2. Turning our attention to the vertices outside of the cycle, we see that there are (${n-2i\choose n/2-i-b+c}$ ways to select the in- and out-vertices that are not in $C$.
We now consider the distribution of special points between our orientations. There are (${7\choose 2}^{i-b}$ ways to select special points in vertices that have an in-degree of 2 in the cycle and are out-vertices in the graph. Likewise, there are (${7\choose 2}^{i-c}$ ways to select special points in vertices that have an out-degree of 2 in the cycle and which are in-vertices in the graph. The special points for vertices with in-degree 2 in $C$ and which are in-vertices and for vertices with out-degree 2 in $C$ and which are out-vertices do not need to be identified. There are $7^{k-2i}$ ways to select special points among vertices in $C$ which have an in-degree of 1 and an out-degree of 1. There are (${9\choose 2}^{n-k}$ ways to select special points for vertices outside of $C$. Given that $9n/2-k$ must be in-points, there are $(9n/2-k)!$ different ways to pair in-points with out-points. We are now ready to verify Lemma 2.5.
@@
We calculate $\frac{\mathbb{E}[YX_{k}]}{\mathbb{E}[Y]}$ as follows:
@@
Since $n\to\infty$ and $k$ is finite, Stirling’s approximation implies that $\frac{(n-2i)!}{n!}\sim\left(\frac{n}{e}\right)^{-2i}$ and
@@
Returning to our calculation of $\frac{\mathbb{E}[YX_{k}]}{\mathbb{E}Y}$ gives us
@@
Since Corollary 2.19 from [4] stipulates that the cycles are asymptotically independent, Lemma 1 by Janson from [12] implies that for every finite sequence $j_{1},j_{2},\ldots,j_{k}$ of non-negative integers, the following holds:
@@
where $\delta_{i}=\frac{\mu_{i}}{\lambda_{i}}-1$ to satisfy condition (c). For the $i$ that we concerned ourselves with above $\delta_{i}=-\left(\frac{2}{9}\right)^{i}$.